### 23 Mar 2026 v1

# Alternating geometric progressions modulo one and Sturmian words

Abstract
Let $b\geq 2$ be an integer. Using Sturmian words we describe all irrational real numbers $\xi$ such that the image in $\mathbb{R}/\mathbb{Z}$ of the sequence $(\xi(-b)^{n})_{n\geq 0}$ is contained in an interval of length $b^{-1}+b^{-2}-b^{-3}$. In previous work we showed that the image cannot be contained in a shorter interval.
1 Introduction
Distribution modulo one of geometric progressions has been the subject of extensive studies. The famous question considered by [Mahler] on the case of common ratio $3/2$ remains open. A theorem of Dubickas [Dubickas-BLMS, Theorem 1] implies that for any integer $b\geq 2$ and any irrational real number $\xi$, the numbers $\{\xi b^{n}\}$, $n\geq 0$ cannot be contained in an interval of length less than $1/b$. Here $\{x\}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}x-\lfloor x\rfloor$ denotes the fractional part of a real number $x$. Bugeaud and Dubickas [BD, Theorem 2.1] showed that for an irrational real number $\xi$, the numbers $\{\xi b^{n}\}$, $n\geq 0$ are contained in an interval of length $1/b$ if and only if the representation $w_{1}w_{2}\dots$ of $\{\xi\}$ in base $b$ is given by a Sturmian word $w_{1}w_{2}\dots$ over an alphabet consisting of two adjacent integers. As pointed out in [AG], the case $b=2$ was proved by Veerman [Veerman, Theorem 2.1] and rediscovered several times.
For alternating geometric progressions with integral common ratio, we showed in previous work [LZ, Theorem 1.1] that for any irrational real number $\xi$, $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is not contained in any interval of $\mathbb{R}/\mathbb{Z}$ of length less than $b^{-1}+b^{-2}-b^{-3}$. Here $\pi\colon\mathbb{R}\to\mathbb{R}$$\mathbb{Z}$ denotes the projection. In this paper, we describe the irrational real numbers $\xi$ such that $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is contained in an interval of $\mathbb{R}/\mathbb{Z}$ of length $b^{-1}+b^{-2}-b^{-3}$.
To state our results we need some terminology from combinatorics on words. An infinite word $w_{1}w_{2}\dots$ is said to be aperiodic if there does not exist integers $k,N\geq 1$ such that $w_{n}=w_{n+k}$ for all $n\geq N$. The slope of an integer-valued infinite word $w_{1}w_{2}\dots$ is defined to be $\lim_{n\to\infty}\frac{\sum_{i=1}^{n}w_{i}}{n}$. An integer-valued infinite word with irrational slope is necessarily aperiodic. An infinite word $s_{1}s_{2}\dots$ over the alphabet $\{0,1\}$ is Sturmian if and only if there exists an irrational $\theta\in(0,1)$ and a $\rho\in\mathbb{R}$ such that
(1.1) $s_{n}=\lfloor(n+1)\theta+\rho\rfloor-\lfloor n\theta+\rho\rfloor$
for all $n\geq 1$ or
(1.2) $s_{n}=\lceil(n+1)\theta+\rho\rceil-\lceil n\theta+\rho\rceil$
for all $n\geq 1$. These words have slope $\theta$. The characteristic Sturmian word $\mathbf{c}_{\theta}=c_{1}c_{2}\dots$ of slope $\theta$ is defined by
$c_{n}=\lfloor(n+1)\theta\rfloor-\lfloor n\theta\rfloor=\lceil(n+1)\theta\rceil-\lceil n\theta\rceil$
for all $n\geq 1$. We refer to [Lothaire, Chapter 2] for a detailed introduction to Sturmian words.
For a word $\mathbf{w}=w_{1}w_{2}\dots$, we write $D\mathbf{w}=w_{1}w_{1}w_{2}w_{2}\dots$ and $T\mathbf{w}=w_{2}w_{3}\dots$. Note that $\mathbf{w}$ has slope $\theta$ if and only if $D\mathbf{w}$ has slope $\theta$. Moreover, $T$ preserves Sturmian words. For a letter $a$ and $l\geq 0$ (finite or infinite), we write $a^{l}$ for the word $aa\dots$ of length $l$.
Notation 1.1.
We let $\mathcal{D}$ denote the collection of infinite words over the alphabet $\{0,1\}$ that are of the form $0^{l}(D\mathbf{s})$ or $1^{l}(D\mathbf{s})$, where $l\geq 0$ and $\mathbf{s}$ is a Sturmian word. We let $\mathcal{C}$ denote the subset of $\mathcal{D}$ consisting of words $\mathbf{w}\in\mathcal{D}$ such that $T^{n}\mathbf{w}$ equals $011(D\mathbf{c})$ or $100(D\mathbf{c})$ for some $n\geq 0$ and some characteristic Sturmian word $\mathbf{c}$ (which is necessarily $\mathbf{c}_{\theta}$, where $\theta$ is the slope of $\mathbf{w}$).
For a word $\mathbf{w}=w_{1}w_{2}\dots$ over a finite set of integers and a real number $r$ satisfying $\lvert r\rvert<1$, we write $t_{r}(\mathbf{w})=\sum_{i=1}^{\infty}w_{i}r^{i}\in\mathbb{R}$. In particular, $t_{-1/b}(\mathbf{w})$ is the real number $w_{1}w_{2}\dots$ in base b-b. For comments on the history of negative-base number systems, see [Knuth, Section 4.1].
Our main result is the following.
Theorem 1.2.
Let $b\geq 2$ be an integer and let $\xi$ be an irrational real number. Then the sequence $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is contained in an interval of $\mathbb{R}/\mathbb{Z}$ of length $b^{-1}+b^{-2}-b^{-3}$ if and only if $\xi=g/(b+1)+t_{-1/b}(\mathbf{w})$, where $g$ is an integer and $\mathbf{w}\in\mathcal{D}$. In this case, $\xi$ is transcendental and the endpoints of the interval are
(1.3) $\pi(g/(b+1)+t_{-1/b}(100D\mathbf{c}_{\theta})),\quad\pi(g/(b+1)+t_{-1/b}(011D\mathbf{c}_{\theta})),$
where $\theta$ denotes the slope of $\mathbf{w}$, and the sequence is contained in a semiopen interval with these endpoints. In particular, the sequence is contained in an open interval of $\mathbb{R}/\mathbb{Z}$ of length $b^{-1}+b^{-2}-b^{-3}$ if and only if $\xi=g/(b+1)+t_{-1/b}(\mathbf{w})$, where $g$ is an integer and $\mathbf{w}\in\mathcal{D}\backslash\mathcal{C}$.
Since $t_{-1/b}$ restricts to an injection $\mathcal{D}\to\mathbb{R}\backslash\mathbb{Q}$, Theorem 1.2 implies that the set of open intervals $I\subseteq\mathbb{R}/\mathbb{Z}$ of length $b^{-1}+b^{-2}-b^{-3}$ such that there exists an irrational real number $\xi$ with $(\pi(\xi(-b)^{n}))_{n\geq 0}$ contained in $I$ has the cardinality of the continuum. Indeed, for any irrational $\theta\in(0,1)$, $D\mathbf{c}_{\theta}\in\mathcal{D}\backslash\mathcal{C}$ and consequently $\xi=t_{-b}(D\mathbf{c}_{\theta})$ satisfies this property with $I=(t_{-1/b}(100D\mathbf{c}_{\theta})),t_{-1/b}(011D\mathbf{c}_{\theta})))$.
The aforementioned result of Veerman [Veerman, Theorem 2.1] implies the following characterization of Sturmian words: An aperiodic word $\mathbf{w}$ over $\{0,1\}$ is Sturmian if and only if there exists an infinite word $\mathbf{u}$ over $\{0,1\}$ such that $0\mathbf{u}\preceq_{\mathrm{lex}}\mathbf{w}\preceq_{\mathrm{lex}}1\mathbf{u}$. Here $\preceq_{\mathrm{lex}}$ denotes the lexicographical order. We will briefly review this characterization in Section 3. For a nice survey of this subject, see [AG].
In negative-base number systems, the order on real numbers is compatible with the alternate order $\preceq_{\mathrm{alt}}$ on words rather than the lexicographical order $\preceq_{\mathrm{lex}}$. For integer-valued infinite words $\mathbf{v}=v_{1}v_{2}\dots$ and $\mathbf{w}=w_{1}w_{2}\dots$, we write $\mathbf{v}\preceq_{\mathrm{alt}}\mathbf{w}$ if $N\mathbf{v}\preceq_{\mathrm{lex}}N\mathbf{w}$, where $N\mathbf{v}=(-v_{1})v_{2}(-v_{3})v_{4}\dots$ and similarly for $N\mathbf{w}$. It is easy to see that, for infinite words $\mathbf{v}$ and $\mathbf{w}$ over $\{0,1\}$ and $0<r<1/2$, $\mathbf{v}\preceq_{\mathrm{alt}}\mathbf{w}$ if and only if $t_{-r}(\mathbf{v})\leq t_{-r}(\mathbf{w})$ (Lemma 4.7). Theorem 1.2 implies the following characterization of the class $\mathcal{D}$.
Corollary 1.3.
An aperiodic word $\mathbf{w}$ over $\{0,1\}$ belongs to $\mathcal{D}$ if and only if there exists an infinite word $\mathbf{u}$ over $\{0,1\}$ such that $100\mathbf{u}\preceq_{\mathrm{alt}}T^{n}\mathbf{w}\preceq_{\mathrm{alt}}011\mathbf{u}$ for all $n\geq 0$. In this case, $\mathbf{u}=D\mathbf{c}_{\theta}$ and
$100\mathbf{u}=\inf\nolimits^{\mathrm{alt}}_{n\geq 0}T^{n}\mathbf{w},\quad 011\mathbf{u}=\sup\nolimits^{\mathrm{alt}}_{n\geq 0}T^{n}\mathbf{w}.$
Here $\theta$ denotes the slope of $\mathbf{w}$, and $\inf^{\mathrm{alt}}$ and $\sup^{\mathrm{alt}}$ denote respectively the infimum and supremum with respect to the order $\preceq_{\mathrm{alt}}$.
More generally, we prove the following result about $t_{-r}$.
Theorem 1.4.
Let $\mathbf{v}=v_{1}v_{2}\dots$ be an aperiodic word over a finite set of integers and let $r$ be a real number satisfying $0<r<1$. Then the sequence $(t_{-r}(T^{n}\mathbf{v}))_{n\geq 0}$ is contained in an interval of length $r+r^{2}-r^{3}$ if and only if there exist $g\in\mathbb{Z}$ and $\mathbf{w}=w_{1}w_{2}\dotso\in\mathcal{D}$ such that $v_{i}=-g+w_{i}$ for all $i\geq 1$. In this case, the endpoints of the interval are
$gr/(1+r)+t_{-r}(100D\mathbf{c}_{\theta}),\quad gr/(1+r)+t_{-r}(011D\mathbf{c}_{\theta}),$
where $\theta$ denotes the slope of $\mathbf{w}$, and the sequence is contained in a semiopen interval with these endpoints. Furthermore, in this case, the sequence is contained in the open interval with these endpoints if and only if $\mathbf{w}\notin\mathcal{C}$.
For alternating geometric progressions with rational common ratio p/q-p/q, where $p>q>1$ are integers, we showed in [LZ, Theorem 1.1] that for any real number $\xi\neq 0$, the sequence $(\pi(\xi(-p/q)^{n}))_{n\geq 0}$ is not contained in any interval of $\mathbb{R}/\mathbb{Z}$ of length less than $(r+r^{2}-r^{3})/q$. In subsequent work [LZ3], we will use Theorem 1.4 to study the question whether the sequence can lie in an interval of length equal to $(r+r^{2}-r^{3})/q$. The analogous question for powers of positive rational numbers was studied by several authors ([FLP], [Bugeaud-linear], [Dubickas-small], [Dubickas-large]).
This paper is organized as follows. In Section 2, we present some preliminary results. In Section 3, we state and prove an analogue of Theorem 1.4 for $t_{r}$ and deduce the theorems of Veerman and Bugeaud–Dubickas. This serves as a warm-up for the more involved proofs, given in Section 4, of the results stated in the introduction.
Acknowledgments
This work was partially supported by National Natural Science Foundation of China (grant numbers 12125107, 12271037, 12288201), Chinese Academy of Sciences Project for Young Scientists in Basic Research (grant number YSBR-033).
2 Preliminaries
The following characterization of Sturmian words as balanced aperiodic words was proved by Morse and Hedlund (see [MH2] or [Lothaire, Theorem 2.1.13]).
Lemma 2.1.
Let $\mathbf{s}$ be an aperiodic word over the alphabet $\{0,1\}$. Then $\mathbf{s}$ is Sturmian if and only if there does not exist any finite word $\mathbf{u}$ such that $0\mathbf{u}0$ and $1\mathbf{u}1$ are subwords of $\mathbf{s}$.
The “only if” part of Lemma 2.1 can be strengthened as follows.
Lemma 2.2.
Let $\mathbf{s}$ and $\mathbf{s}^{\prime}$ be Sturmian words over $\{0,1\}$ of the same slope. Then there does not exist any finite word $\mathbf{u}$ such that $0\mathbf{u}0$ is a subword of $\mathbf{s}$ and $1\mathbf{u}1$ is a subword of $\mathbf{s}^{\prime}$.
Proof.
This follows immediately from the “only if” part of Lemma 2.1 and the fact that $\mathbf{s}$ and $\mathbf{s}^{\prime}$ have the same set of finite subwords [Lothaire, Proposition 2.1.18(i)]. More directly, it follows from Lemma 2.3 below. ∎
Lemma 2.3.
Let $s_{1}s_{2}\dots$ be a Sturmian word over $\{0,1\}$ of slope $\theta$. Then $\lvert\theta k-\sum_{i={n}}^{n+k-1}s_{i}\rvert<1$ for all $n,k\geq 1$.
Proof.
This follows from an easy computation. In the case (1.1), we have
$\theta k-\sum_{i={n}}^{n+k-1}s_{i}=\theta k-(\lfloor(n+k)\theta+\rho\rfloor-\lfloor n\theta+\rho\rfloor)=\{(n+k)\theta+\rho\}-\{n\theta+\rho\}\in(-1,1).$
The same holds in the case (1.2) after replacing $\lfloor\cdot\rfloor$ by $\lceil\cdot\rceil$ and $\{\cdot\}$ by $\{-(\cdot)\}$. ∎
Lemma 2.4.
Let $\theta\in(0,1)$ be irrational. Then $0\mathbf{c}_{\theta}$ and $1\mathbf{c}_{\theta}$ are Sturmian.
Proof.
This is clear by taking $\rho=-\theta$ in (1.1) and (1.2). ∎
Lemma 2.5.
Let $0<r<1/2$. Then $t_{r}$ restricts to an order-preserving injection $(\{0,1\}^{\mathbb{N}},\preceq_{\mathrm{lex}})\to(\mathbb{R},\leq)$.
Proof.
Let $\mathbf{s}=s_{1}s_{2}\dots$, $\mathbf{s}^{\prime}=s^{\prime}_{1}s^{\prime}_{2}\dots\in\{0,1\}^{\mathbb{N}}$ with $\mathbf{s}\prec_{\mathrm{lex}}\mathbf{s}^{\prime}$. Then there exists $n\geq 0$ such that $s_{i}=s^{\prime}_{i}$ for all $i<n$ and $s_{n}<s^{\prime}_{n}$. Thus
$t_{r}(\mathbf{s}^{\prime})-t_{r}(\mathbf{s})=\sum_{i=1}^{\infty}(s^{\prime}_{i}-s_{i})r^{i}\geq r^{n}-\sum_{i=n+1}^{\infty}r^{i}=r^{n}(1-\frac{r}{1-r})>0.$
Lemma 2.6.
Let $0<r<1$ and let $\mathbf{s}$ and $\mathbf{s}^{\prime}$ be Sturmian words over $\{0,1\}$ of the same slope. Then $\lvert t_{r}(\mathbf{s}^{\prime})-t_{r}(\mathbf{s})\rvert\leq r$. Moreover, if $\mathbf{s}\prec_{\mathrm{lex}}\mathbf{s}^{\prime}$, then $t_{r}(\mathbf{s})<t_{r}(\mathbf{s}^{\prime})$.
Proof.
We may assume $\mathbf{s}\prec_{\mathrm{lex}}\mathbf{s}^{\prime}$. Write $\mathbf{s}=s_{1}s_{2}\dots$ and $\mathbf{s}^{\prime}=s^{\prime}_{1}s^{\prime}_{2}\dots$. Then
$t_{r}(\mathbf{s}^{\prime})-t_{r}(\mathbf{s})=\sum_{n=1}^{\infty}(s^{\prime}_{n}-s_{n})r^{n}.$
By Lemma 2.2, the sequence $(s^{\prime}_{n}-s_{n})_{n\geq 1}$ with values in $\{-1,0,1\}$ alternates in sign after removing all zeroes. Moreover, the leading nonzero term is positive by the assumption $\mathbf{s}\prec_{\mathrm{lex}}\mathbf{s}^{\prime}$. Thus $0<t_{r}(\mathbf{s}^{\prime})-t_{r}(\mathbf{s})\leq r$. ∎
The notation $t_{r}$ extends naturally to finite words. Let $\Sigma$ be a finite set of integers. For a finite word $\mathbf{u}=u_{1}\dots u_{k}$ over $\Sigma$, we write $t_{r}(\mathbf{u})=\sum_{i=1}^{k}u_{i}r^{i}$. Then, for any finite or infinite word $\mathbf{v}$ over $\Sigma$, we have
$t_{r}(\mathbf{u}\mathbf{v})-r^{k}t_{r}(\mathbf{v})=t_{r}(\mathbf{u}).$
In particular, for any infinite word $\mathbf{w}$ over $\Sigma$, if $T^{n}\mathbf{w}=\mathbf{u}T^{n+k}\mathbf{w}$ and $T^{m}\mathbf{w}=\mathbf{u}^{\prime}T^{m+k}\mathbf{w}$, where $\mathbf{u}$ and $\mathbf{u}^{\prime}$ are words of length $k$, then
(2.1) $(t_{r}(T^{n}\mathbf{w})-t_{r}(T^{m}\mathbf{w}))-r^{k}(t_{r}(T^{n+k}\mathbf{w})-t_{r}(T^{m+k}\mathbf{w}))=t_{r}(\mathbf{u})-t_{r}(\mathbf{u}^{\prime}).$
3 The positive case
As mentioned in the introduction, the following analogue of Theorem 1.4 was proved by Veerman [Veerman, Theorem 2.1] in the case $r=1/2$ and by Bugeaud and Dubickas [BD, Theorem 2.1] in the case $r=1/b$, where $b\geq 2$ is an integer.
Proposition 3.1.
Let $\mathbf{v}=v_{1}v_{2}\dots$ be an aperiodic word over a finite set of integers and let $r$ be a real number such that $0<r<1$. Then the sequence $(t_{r}(T^{n}\mathbf{v}))_{n\geq 0}$ is contained in an interval of length $r$ if and only if there exist an integer $g$ and a Sturmian word $\mathbf{w}=w_{1}w_{2}\dots$ over the alphabet $\{0,1\}$ such that $v_{i}=g+w_{i}$ for all $i\geq 1$. In this case, the endpoints of the interval are
$gr/(1-r)+t_{r}(0\mathbf{c}_{\theta}),\quad gr/(1-r)+t_{r}(1\mathbf{c}_{\theta}),$
where $\theta$ denotes the slope of $\mathbf{w}$, and the sequence is contained in a semiopen interval with these endpoints. Furthermore, in this case, the sequence is contained in the open interval with these endpoints if and only if none of the words $T^{n}\mathbf{w}$, $n\geq 1$ is characteristic Sturmian.
Veerman [Veerman, Remark following Lemma 2.4] already noted that (in the case $\mathbf{v}$ over $\{0,1\}$) the first assertion of Proposition 3.1 holds for $0<r<1$ and the second assertion holds for $0<r\leq 1/2$. We will see that the restriction to $r\leq 1/2$ is unnecessary.
Proof.
Assume that $(t_{r}(T^{n}\mathbf{v}))_{n\geq 0}$ is contained in an interval $[A,B]$ of length $r$. Let $h=\max_{n\geq 1}v_{n}$ and $g=\min_{n\geq 1}v_{n}$. Since $\mathbf{v}$ is aperiodic, we have $h>g$. By (2.1),
$2r>(1+r)(B-A)\geq t_{r}(h)-t_{r}(g)=(h-g)r.$
Thus hg=1h-g=1 and consequently $v_{n}\in\{g,g+1\}$ for all $n\geq 1$. Let $w_{n}=v_{n}-g\in\{0,1\}$ and $\mathbf{w}=w_{1}w_{2}\dots$. Then $t_{r}(\mathbf{w})=t_{r}(\mathbf{v})-t_{r}(g^{\infty})=t_{r}(\mathbf{v})-gr/(1-r)$. Thus, up to replacing $\mathbf{v}$ by $\mathbf{w}$ we may assume that $g=0$. Assume that $\mathbf{w}$ is not Sturmian. By Lemma 2.1, since $\mathbf{w}$ is aperiodic but Sturmian, there exists a word $\mathbf{u}$ of length $k\geq 0$ such that $0\mathbf{u}0$ and $1\mathbf{u}1$ are subwords of $\mathbf{w}$. Then, by (2.1), we have
$r+r^{k+3}=(1+r^{k+2})(B-A)\geq t_{r}(1\mathbf{u}1)-t_{r}(0\mathbf{u}0)=r+r^{k+2}.$
Contradiction.
For the other statements of the proposition, we may again assume $g=0$. Let $\mathbf{w}$ be a Sturmian word over $\{0,1\}$ of slope $\theta$ and let $W=\{t_{r}(T^{n}\mathbf{w})\mid n\geq 0\}$. Note that $t_{r}(1\mathbf{c}_{\theta})=t_{r}(0\mathbf{c}_{\theta})+r$. By Lemma 2.6 applied to the Sturmian words $T^{n}\mathbf{w}$, $0\mathbf{c}_{\theta}$, and $1\mathbf{c}_{\theta}$, we have
$t_{r}(T^{n}\mathbf{w})=t_{r}(1\mathbf{c}_{\theta})-(t_{r}(1\mathbf{c}_{\theta})-t_{r}(T^{n}\mathbf{w}))\geq t_{r}(1\mathbf{c}_{\theta})-r=t_{r}(0\mathbf{c}_{\theta}),$
$t_{r}(T^{n}\mathbf{w})=t_{r}(0\mathbf{c}_{\theta})+(t_{r}(T^{n}\mathbf{w})-t_{r}(0\mathbf{c}_{\theta}))\leq t_{r}(0\mathbf{c}_{\theta})+r=t_{r}(1\mathbf{c}_{\theta}).$
Thus $W$ is contained in the interval $t_{r}(0\mathbf{c}_{\theta}),t_{r}(1\mathbf{c}_{\theta})]$ of length $r$. By [LZ, Proposition 2.3], $W$ is not contained in any shorter interval.
By Lemma 2.6, $t_{r}(0\mathbf{c}_{\theta})\in W$ if and only if $T^{n}\mathbf{w}=0\mathbf{c}_{\theta}$ for some $n\geq 0$, and $t_{r}(1\mathbf{c}_{\theta})\in W$ if and only if $T^{m}\mathbf{w}=1\mathbf{c}_{\theta}$ for some $m\geq 0$. If both conditions are satisfied, then $T^{m+1}\mathbf{w}=\mathbf{c}_{\theta}=T^{n+1}\mathbf{w}$ for $m\neq n$, which contradicts the aperiodicity of $\mathbf{w}$. Thus $W$ is contained in $t_{r}(0\mathbf{c}_{\theta}),t_{r}(1\mathbf{c}_{\theta}))$ or $(t_{r}(0\mathbf{c}_{\theta}),t_{r}(1\mathbf{c}_{\theta})]$. Furthermore, $W$ is contained in $(t_{r}(0\mathbf{c}_{\theta}),t_{r}(1\mathbf{c}_{\theta}))$ if and only if $T^{n}\mathbf{w}=\mathbf{c}_{\theta}$ for some $n\geq 1$. ∎
The following statement is taken from [AG, Theorem 1].
Corollary 3.2 (Veerman).
An aperiodic word $\mathbf{w}$ over $\{0,1\}$ is Sturmian if and only if there exists an infinite word $\mathbf{u}$ over $\{0,1\}$ such that $0\mathbf{u}\preceq_{\mathrm{lex}}T^{n}\mathbf{w}\preceq_{\mathrm{lex}}1\mathbf{u}$ for all $n\geq 0$. In this case, $\mathbf{u}=\mathbf{c}_{\theta}$ and
$0\mathbf{u}=\inf\nolimits^{\mathrm{lex}}_{n\geq 0}T^{n}\mathbf{w},\quad 1\mathbf{u}=\sup\nolimits^{\mathrm{lex}}_{n\geq 0}T^{n}\mathbf{w}.$
Here $\theta$ denotes the slope of $\mathbf{w}$, and $\inf^{\mathrm{lex}}$ and $\sup^{\mathrm{lex}}$ denote respectively the infimum and supremum with respect to the order $\preceq_{\mathrm{alt}}$.
Proof.
Take $0<r<1/2$. The corollary follows easily from Proposition 3.1 and Lemma 2.5. In more details, if there exists $\mathbf{u}$ such that $0\mathbf{u}\preceq_{\mathrm{lex}}T^{n}\mathbf{w}\preceq_{\mathrm{lex}}1\mathbf{u}$, then, by Lemma 2.5, $(t_{r}(T^{n}\mathbf{w}))_{n\geq 0}$ is contained in the interval $t_{r}(0\mathbf{u}),t_{r}(1\mathbf{u})]$ of length $r$, which implies that $\mathbf{w}$ is Sturmian by Proposition 3.1. Conversely, assume that $\mathbf{w}$ is Sturmian of slope $\theta$. Then, by Proposition 3.1, for all $n\geq 0$, $t_{r}(0\mathbf{c}_{\theta})\leq t_{r}(T^{n}\mathbf{w})\leq t_{r}(0\mathbf{c}_{\theta})$, which implies $0\mathbf{c}_{\theta}\preceq_{\mathrm{lex}}T^{n}\mathbf{w}\preceq_{\mathrm{lex}}1\mathbf{c}_{\theta}$. If there exists an infinite word $\mathbf{v}$ over $\{0,1\}$ such that $0\mathbf{c}_{\theta}\prec_{\mathrm{lex}}\mathbf{v}\preceq_{\mathrm{lex}}T^{n}\mathbf{w}$ for all $n\geq 0$, then, by Lemma 2.5, $(t_{r}(T^{n}\mathbf{w}))_{n\geq 0}$ is contained in the interval $t_{r}(\mathbf{v}),t_{r}(1\mathbf{c}_{\theta})]$ of length <r<r, which contradicts [LZ, Proposition 2.3]. Thus $0\mathbf{c}_{\theta}=\inf\nolimits^{\mathrm{lex}}_{n\geq 0}T^{n}\mathbf{w}$ and similarly $1\mathbf{c}_{\theta}=\sup\nolimits^{\mathrm{lex}}_{n\geq 0}T^{n}\mathbf{w}$. ∎
Corollary 3.3 (Bugeaud–Dubickas).
Let $b\geq 2$ be an integer and let $\xi$ be an irrational real number. Then the sequence $(\pi(\xi b^{n}))_{n\geq 0}$ is contained in an interval of $\mathbb{R}/\mathbb{Z}$ of length $1/b$ if and only if $\xi=g/(b-1)+t_{1/b}(\mathbf{w})$, where $g$ is an integer and $\mathbf{w}$ is a Sturmian word over $\{0,1\}$. In this case, $\xi$ is transcendental and the endpoints of the interval are
$\pi(g/(b-1)+t_{1/b}(\mathbf{c}_{\theta})),\quad\pi(g/(b-1)+t_{1/b}(\mathbf{c}_{\theta})),$
where $\theta$ denotes the slope of $\mathbf{w}$, and the sequence is contained in a semiopen interval with these endpoints. Furthermore, in this case, the sequence is contained in the open interval with these endpoints if and only if none of the words $T^{n}\mathbf{w}$, $n\geq 1$ is characteristic Sturmian.
Proof.
This is essentially [BD, Theorem 2.1], except that here we do not a priori exclude the case where $\pi(0)$ is in the interior of the interval or is the upper endpoint. Assume that $(\pi(\xi b^{n}))_{n\geq 0}$ is contained in an interval $I\subseteq\mathbb{R}/\mathbb{Z}$ of length $1/b$. Choose $\eta\in\mathbb{R}$ such that $\pi(\eta)\notin I$. Then $I$ lifts to a unique interval $\tilde{I}\subseteq(\eta,\eta+1)$. For $n\geq 0$, let $y_{n}=\{\xi b^{n}-\eta\}+\eta\in(\eta,\eta+1)$ and $v_{n}=by_{n}-y_{n+1}\in\mathbb{Z}\cap((b-1)\eta-1,(b-1)\eta+b)$. Since $\pi(y_{n})=\pi(\xi b^{n})$, we have $y_{n}\in\tilde{I}$ for all $n\geq 0$. Let $\mathbf{v}=v_{0}v_{1}\dots$. Then $t_{1/b}(T^{n}\mathbf{v})=\sum_{i=0}^{\infty}v_{n+i}b^{-(i+1)}=y_{n}$. Since $y_{0}$ is irrational, $\mathbf{v}$ is aperiodic. Then, by Proposition 3.1, there exists $g\in\mathbb{Z}$ and a Sturmian word $\mathbf{w}=w_{0}w_{1}\dots$ over $\{0,1\}$ such that $v_{n}=g+w_{n}$ for all $n\geq 0$. Thus $\pi(\xi)=\pi(y_{0})=\pi(g/(b-1)+t_{1/b}(\mathbf{w}))$. By a theorem of Ferenczi and Mauduit [FM, Proposition 2], $t_{1/b}(\mathbf{w})$ is transcendental and consequently the same holds for $\xi$. The other assertions of Corollary 3.3 follow easily from the corresponding assertions of Proposition 3.1 and the formula $\pi(t_{1/b}(T^{n}\mathbf{v}))=\pi(b^{n}t_{1/b}(\mathbf{v}))$. ∎
4 Proofs of the main results
We need a slight generalization of Sturmian words.
Notation 4.1.
For $l\geq 0$, we let $\mathcal{S}_{l}$ denote the collection of infinite words over $\{0,1\}$ that are of the form $0^{l}\mathbf{s}$ or $1^{l}\mathbf{s}$, where $\mathbf{s}$ is a Sturmian word. We write $\mathcal{S}=\bigcup_{l\geq 0}\mathcal{S}_{l}$.
The following characterization of the class $\mathcal{S}$ parallels Lemma 2.1.
Lemma 4.2.
Let $\mathbf{w}$ be an aperiodic word over $\{0,1\}$. Then $\mathbf{w}\in\mathcal{S}$ if and only there does not exist a finite word $\mathbf{u}$ such that $01\mathbf{u}1$ and $10\mathbf{u}0$ are subwords of $\mathbf{w}$.
Proof.
The “only if” part. We have $\mathbf{w}=0^{l}\mathbf{s}$ or $\mathbf{w}=1^{l}\mathbf{s}$ with $l\geq 0$ and $\mathbf{s}$ Sturmian. If $01\mathbf{u}1$ and $10\mathbf{u}0$ are subwords of $\mathbf{w}$, then $0\mathbf{u}0$ and $1\mathbf{u}1$ are subwords of $\mathbf{s}$, which contradicts Lemma 2.1.
The “if” part. We have $\mathbf{w}=0^{l}\mathbf{s}$ with $l\geq 1$ and $\mathbf{s}=1\dots$ or $\mathbf{w}=1^{l}\mathbf{s}$ with $l\geq 1$ and $\mathbf{s}=0\dots$. Assume that $\mathbf{s}$ is not Sturmian. By Lemma 2.1, there exists a finite word $\mathbf{u}$ such that $0\mathbf{u}0$ and $1\mathbf{u}1$ are subwords of $\mathbf{s}$. Assume that $01\mathbf{u}1$ is a subword of $\mathbf{w}$. Then, by assumption, $10\mathbf{u}0$ is not a subword of $\mathbf{w}$. It follows that $00\mathbf{u}0$ is a subword of $\mathbf{s}$. Let $k\geq 0$ be the number of leading 0s in $\mathbf{u}$. Then $010^{k}1$ is a subword of $01\mathbf{u}1$ and thus a subword of $\mathbf{w}$. Moreover, $0^{k+2}$ is a subword of $00\mathbf{u}$ and thus a subword of $\mathbf{s}$, and consequently $10^{k+2}=100^{k}0$ is a subword of $\mathbf{w}$. This contradicts the assumption. Therefore, $01\mathbf{u}1$ is not a subword of $\mathbf{w}$ and consequently $11\mathbf{u}1$ is a subword of $\mathbf{s}$. Similarly, $00\mathbf{u}0$ is a subword of $\mathbf{s}$. It follows that $011$ and $100$ are subwords of $\mathbf{w}$. Contradiction. ∎
The order-preserving property in Lemma 2.6 extends to $\mathcal{S}$.
Lemma 4.3.
Let $0<r<1$. Let $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{S}$ be of the same slope and such that $\mathbf{s}\prec_{\mathrm{lex}}\mathbf{s}^{\prime}$. Then $t_{r}(\mathbf{s})<t_{r}(\mathbf{s}^{\prime})$.
Proof.
Assume $\mathbf{s}\in\mathcal{S}_{l}$ and $\mathbf{s}^{\prime}\in\mathcal{S}_{l^{\prime}}$. Let $n=\max(l,l^{\prime})$. Write $\mathbf{s}=\mathbf{u}T^{n}\mathbf{s}$ and $\mathbf{s}^{\prime}=\mathbf{u}^{\prime}T^{n}\mathbf{s}^{\prime}$, where $\mathbf{u}$ and $\mathbf{u}^{\prime}$ are words of length $n$. Then $\{\mathbf{u},\mathbf{u}^{\prime}\}\cap\{0^{n},1^{n}\}$ is nonempty. Moreover, $T^{n}\mathbf{s}$ and $T^{n}\mathbf{s}^{\prime}$ are Sturmian. We have
$t_{r}(\mathbf{s}^{\prime})-t_{r}(\mathbf{s})=(t_{r}(\mathbf{u}^{\prime})-t_{r}(\mathbf{u}))+r^{n}(t_{r}(T^{n}\mathbf{s}^{\prime})-t_{r}(T^{n}\mathbf{s}))$
There are two cases for $\mathbf{s}\prec_{\mathrm{lex}}\mathbf{s}^{\prime}$.
Case $\mathbf{u}=\mathbf{u}^{\prime}$ and $T^{n}\mathbf{s}\prec_{\mathrm{lex}}T^{n}\mathbf{s}^{\prime}$. In this case, $t_{r}(\mathbf{s}^{\prime})-t_{r}(\mathbf{s})=r^{n}(t_{r}(T^{n}\mathbf{s}^{\prime})-t_{r}(T^{n}\mathbf{s}))>0$ by Lemma 2.6.
Case $\mathbf{u}\prec_{\mathrm{lex}}\mathbf{u}^{\prime}$. Since either $\mathbf{u}=0^{n}$ or $\mathbf{u}^{\prime}=1^{n}$, we have $t_{r}(\mathbf{u}^{\prime})-t_{r}(\mathbf{u})\geq r^{n}$. Moreover, $\lvert t_{r}(T^{n}\mathbf{s}^{\prime})-t_{r}(T^{n}\mathbf{s})\rvert\leq r$ by Lemma 2.6. Thus $t_{r}(\mathbf{s}^{\prime})-t_{r}(\mathbf{s})\geq r^{n}-r^{n+1}>0$. ∎
The classes $\mathcal{S}$ and $\mathcal{D}$ are related as follows.
Lemma 4.4.
We have $\mathcal{D}=D(\mathcal{S})\coprod T(D(\mathcal{S}))$.
Proof.
By definition $D(\mathcal{S})\subseteq\mathcal{D}$ and $T(D(\mathcal{S}_{l}))\subseteq\mathcal{D}$ for $l\geq 1$. For $\mathbf{s}=s_{1}\dots$ Sturmian, $T(D(\mathbf{s}))=s_{1}D(T(\mathbf{s}))\in\mathcal{D}$. Thus $D(\mathcal{S})\cup T(D(\mathcal{S}))\subseteq\mathcal{D}$.
For $\mathbf{w}\in\mathcal{D}$, we have $\mathbf{w}=a^{l}D\mathbf{s}$ for $a\in\{0,1\}$, $l\geq 0$ and $\mathbf{s}$ Sturmian. If $l=2k$ is even, then $\mathbf{w}=D(0^{k}\mathbf{s})$. If l=2k1l=2k-1 is odd, then $\mathbf{w}=T(D(0^{k}\mathbf{s}))$. Thus $\mathcal{D}=D(\mathcal{S})\cup T(D(\mathcal{S}))$.
Finally, for $\Sigma=\{0,1\}$, $D(\Sigma^{\mathbb{N}})\cap T(D(\Sigma^{\mathbb{N}}))=\{0^{\infty},1^{\infty}\}$. Thus $D(\mathcal{S})\cap T(D(\mathcal{S}))$ is empty. ∎
Since $\mathcal{S}$ is stable under $T$, it follows from Lemma 4.4 (or the definition) that $\mathcal{D}$ is also stable under $T$.
Next we analyze the alternate order on $\mathcal{D}$. It is convenient to do this for a more general class of words.
Notation 4.5.
Let $\Sigma=\{0,1\}$, $\mathcal{P}=D(\Sigma^{\mathbb{N}})\cup T(D(\Sigma^{\mathbb{N}}))$. We put
$\mathcal{P}_{1^{\infty}]}$ $=\{\mathbf{w}\in\mathcal{P}\mid\mathbf{w}\preceq_{\mathrm{alt}}1^{\infty}\},$
$\mathcal{P}_{[1^{\infty},0^{\infty}]}$ $=\{\mathbf{w}\in\mathcal{P}\mid 1^{\infty}\preceq_{\mathrm{alt}}\mathbf{w}\preceq_{\mathrm{alt}}0^{\infty}\},$
$\mathcal{P}_{[0^{\infty}}$ $=\{\mathbf{w}\in\mathcal{P}\mid 0^{\infty}\preceq_{\mathrm{alt}}\mathbf{w}\}.$
Lemma 4.6.
  1. 1.
    The map $(\Sigma^{\mathbb{N}},\preceq_{\mathrm{lex}})\to(\mathcal{P}_{1^{\infty}]},\preceq_{\mathrm{alt}})$ carrying $\mathbf{u}$ to $1D\mathbf{u}$ is an order-preserving bijection.
  2. 2.
    The map $(\Sigma^{\mathbb{N}},\preceq_{\mathrm{lex}})\to(\mathcal{P}_{[1^{\infty},0^{\infty}]},\preceq_{\mathrm{alt}})$ carrying $\mathbf{u}$ to $D\mathbf{u}$ is an order-reversing bijection.
  3. 3.
    The map $(\Sigma^{\mathbb{N}},\preceq_{\mathrm{lex}})\to(\mathcal{P}_{[0^{\infty}},\preceq_{\mathrm{alt}})$ carrying $\mathbf{u}$ to $0D\mathbf{u}$ is an order-preserving bijection.
Proof.
The bijectivity follows from the fact that for any $\mathbf{w}\in\Sigma^{\mathbb{N}}$, $\mathbf{w}\prec_{\mathrm{alt}}1^{\infty}$ if and only if the number of leading $1$s in $\mathbf{w}$ is odd and $0^{\infty}\prec_{\mathrm{alt}}\mathbf{w}$ if and only if the number of leading 0s in $\mathbf{w}$ is odd. The preservation and reversal of the order follow from the definition. ∎
Lemma 4.7.
Let $0<r<1/2$. Then $t_{-r}$ restricts to an order-preserving injection $(\Sigma^{\mathbb{N}},\preceq_{\mathrm{alt}})\to(\mathbb{R},\leq)$.
Proof.
Let $\mathbf{s}=s_{1}s_{2}\dots$, $\mathbf{s}^{\prime}=s^{\prime}_{1}s^{\prime}_{2}\dots\in\{0,1\}^{\mathbb{N}}$ with $\mathbf{s}\prec_{\mathrm{alt}}\mathbf{s}^{\prime}$. Then there exists $n\geq 0$ such that $s_{i}=s^{\prime}_{i}$ for all $i<n$ and $(-1)^{n}(s^{\prime}_{n}-s_{n})=1$. Thus
$t_{-r}(\mathbf{s}^{\prime})-t_{-r}(\mathbf{s})=\sum_{i=1}^{\infty}(s^{\prime}_{i}-s_{i})(-r)^{i}\geq r^{n}-\sum_{i=n+1}^{\infty}r^{i}=r^{n}(1-\frac{r}{1-r})>0.$
Lemma 4.8.
Let $0<r<1$ and let $\mathbf{w}\in\mathcal{P}$.
  1. 1.
    If $\mathbf{w}\preceq_{\mathrm{alt}}1^{\infty}$, then $t_{-r}(\mathbf{w})\leq t_{-r}(1^{\infty})$ and equality holds if and only if $\mathbf{w}=1^{\infty}$.
  2. 2.
    If $1^{\infty}\preceq_{\mathrm{alt}}\mathbf{w}\preceq_{\mathrm{alt}}0^{\infty}$, then $t_{-r}(1^{\infty})\leq t_{-r}(\mathbf{w})\leq t_{-r}(0^{\infty})$ and the first (resp. second) equality holds if and only if $\mathbf{w}=1^{\infty}$ (resp. $\mathbf{w}=0^{\infty}$).
  3. 3.
    If $0^{\infty}\preceq_{\mathrm{alt}}\mathbf{w}$, then $t_{-r}(0^{\infty})\leq t_{-r}(\mathbf{w})$ and equality holds if and only if $\mathbf{w}=0^{\infty}$.
Proof.
The inequalities follow from Lemma 4.6 and the formulas
(4.1) $t_{-r}(D\mathbf{u})=-(r^{-1}-1)t_{r^{2}}(\mathbf{u}),\quad t_{-r}(aD\mathbf{u})=-ar+(1-r)t_{r^{2}}(\mathbf{u}),$
and $t_{r^{2}}(\mathbf{u})\in[t_{r^{2}}(0^{\infty}),t_{r^{2}}(1^{\infty})]$, where $a\in\Sigma$ and $\mathbf{u}\in\Sigma^{\mathbb{N}}$. The equality conditions follow from the fact that $t_{r^{2}}(\mathbf{u})=t_{r^{2}}(a^{\infty})$ if and only if $\mathbf{u}=a^{\infty}$. ∎
Lemma 4.9.
Let $\mathbf{w}\in\mathcal{D}$ and let $\theta$ be the slope of $\mathbf{w}$. Then $100D\mathbf{c}_{\theta}\preceq_{\mathrm{alt}}\mathbf{w}\preceq_{\mathrm{alt}}011D\mathbf{c}_{\theta}$.
Proof.
Assume $\mathbf{w}\preceq_{\mathrm{alt}}100D\mathbf{c}_{\theta}$. Then $\mathbf{w}=10\dots$. Thus $\mathbf{w}=1D\mathbf{s}$, where $\mathbf{s}$ is Sturmian of slope $\theta$. Since $\mathbf{w}\preceq_{\mathrm{alt}}1D(0\mathbf{c}_{\theta})$, we have $\mathbf{s}\preceq_{\mathrm{lex}}0\mathbf{c}_{\theta}$ by Lemma 4.6. By Corollary 3.2, we must have $\mathbf{s}=0\mathbf{c}_{\theta}$ and consequently $\mathbf{w}=100D\mathbf{c}_{\theta}$. Similarly, $011D\mathbf{c}_{\theta}\preceq_{\mathrm{alt}}\mathbf{w}$ implies $\mathbf{w}=011D\mathbf{c}_{\theta}$. ∎
Lemma 4.10.
Let $0<r<1$. Let $\mathbf{w},\mathbf{w}^{\prime}\in\mathcal{D}$ be of the same slope and such that $\mathbf{w}\prec_{\mathrm{alt}}\mathbf{w}^{\prime}$. Then $t_{-r}(\mathbf{w})<t_{-r}(\mathbf{w}^{\prime})$.
Proof.
By Lemma 4.8, we may assume that one of the three intervals in Notation 4.5 contain both $\mathbf{w}$ and $\mathbf{w}^{\prime}$.
Case $\mathbf{w},\mathbf{w}^{\prime}\in\mathcal{P}_{1^{\infty}]}$ or $\mathbf{w},\mathbf{w}^{\prime}\in\mathcal{P}_{[0^{\infty}}$. By Lemma 4.6, $\mathbf{w}=aD\mathbf{u}$, $\mathbf{w}^{\prime}=aD\mathbf{u}^{\prime}$, where $a\in\{0,1\}$ and $\mathbf{u}\prec_{\mathrm{lex}}\mathbf{u}^{\prime}$. Then $t_{r^{2}}(\mathbf{u})<t_{r^{2}}(\mathbf{u}^{\prime})$ by Lemma 2.6 and consequently $t_{-r}(\mathbf{w})<t_{-r}(\mathbf{w}^{\prime})$ by (4.1).
Case $\mathbf{w},\mathbf{w}^{\prime}\in\mathcal{P}_{[1^{\infty},0^{\infty}]}$. By Lemma 4.6, $\mathbf{w}=D\mathbf{u}$, $\mathbf{w}^{\prime}=D\mathbf{u}^{\prime}$, where $\mathbf{u}^{\prime}\prec_{\mathrm{lex}}\mathbf{u}$. Then $t_{r^{2}}(\mathbf{u}^{\prime})<t_{r^{2}}(\mathbf{u})$ by Lemma 2.6 and consequently $t_{-r}(\mathbf{w})<t_{-r}(\mathbf{w}^{\prime})$ by (4.1). ∎
Proof of Theorem 1.4.
The beginning of the proof of the “only if” part of the first assertion is similar to that of [LZ, Theorem 1.2]. Assume that $(t_{-r}(T^{n}\mathbf{v}))_{n\geq 0}$ is contained in an interval $[A,B]$ of length $r+r^{2}-r^{3}$. Let $\mu=\max_{n\geq 1}v_{n}$ and $\nu=\min_{n\geq 1}v_{n}$. Since $\mathbf{v}$ is aperiodic, we have $\mu>\nu$. Moreover, there exists a subword $\mu\mu^{\prime}$ of $\mathbf{v}$ with $\mu^{\prime}<\mu$. Similarly, there exists a subword $\nu\nu^{\prime}$ of $\mathbf{v}$ with $\nu^{\prime}>\nu$. If $\mu-\nu\geq 2$, then, by (2.1),
$2r-r(1-r)(1-r^{3})=(1+r^{2})(B-A)\geq t_{-r}(\nu\nu^{\prime})-t_{-r}(\mu\mu^{\prime})\\ =-(\nu-\mu)r+(\nu^{\prime}-\mu^{\prime})r^{2}\geq(\mu-\nu)r-(\mu-\nu-2)r^{2}\geq 2r,$
which is a contradiction. Thus $\mu-\nu=1$ and consequently $v_{n}\in\{\nu,\nu+1\}$ for all $n\geq 1$. Let $g=-\nu$, $w_{n}=v_{n}+g\in\{0,1\}$ and $\mathbf{w}=w_{1}w_{2}\dots$. Then $t_{r}(\mathbf{w})=t_{r}(\mathbf{v})+t_{r}(g^{\infty})=t_{r}(\mathbf{v})-gr/(1+r)$. Up to replacing $\mathbf{v}$ by $\mathbf{w}$, we may assume $g=0$. By the above $01$ and $10$ are subwords of $\mathbf{w}$.
Note that $\mathbf{w}$ has a subword of the form $10a$, where $a\in\{0,1\}$. If $010$ is a subword of $\mathbf{w}$, then, by (2.1), we have
$r+r^{2}-r^{3}(1-r)(1-r^{2})=(1+r^{3})(B-A)\geq t_{-r}(010)-t_{-r}(10a)=r+r^{2}+ar^{3}\geq r+r^{2},$
which is a contradiction. Thus $010$ is not a subword of $\mathbf{w}$. Similarly, $101$ is not a subword of $\mathbf{w}$.
Next note that $0111$ and $1000$ are not both subwords of $\mathbf{w}$. Indeed, otherwise, by (2.1), we would have
$r+r^{2}-r^{3}+r^{4}-r^{4}(1-r)(1-r^{2})=(1+r^{4})(B-A)\geq t_{-r}(0111)-t_{-r}(1000)=r+r^{2}-r^{3}+r^{4}.$
Up to replacing $w_{n}$ by $1-w_{n}$ for all $n\geq 1$, we may assume that $0111$ is not a subword of $\mathbf{w}$. Then all blocks of 0s in $\mathbf{w}$ are separated by $11$. Thus $\mathbf{w}=1^{y}0^{z_{0}}110^{z_{1}}110^{z_{2}}\dots$, where $y\geq 0$, $z_{0}\geq 1$, and $z_{n}\geq 2$ for all $n\geq 1$.
Assume that $z_{n}$ is odd for some $n\geq 1$. Then $0110^{z_{n}-1}$ and $10^{z_{n}}1$ are subwords of $\mathbf{w}$. Thus, by (2.1), we have
(4.2) $r+r^{2}-r^{3}+r^{z_{n}+2}-r^{z_{n}+2}(1-r)(1-r^{2})=(1+r^{z_{n}+2})(B-A)\\ \geq t_{-r}(0110^{z_{n}-1})-t_{-r}(10^{z_{n}}1)=r+r^{2}-r^{3}+r^{z_{n}+2}.$
Contradiction. It follows that $z_{n}$ is even for all $n\geq 1$.
The proof now diverges from the proof of [LZ, Theorem 1.2]. We claim that $\mathbf{w}\in\mathcal{P}$. This holds if $y=0$ or $z_{0}$ is even. Assume that $y>0$ and $z_{0}$ is odd. If $z_{0}\leq z_{m}+1$ for some $m\geq 1$, then $0110^{z_{0}-1}$ and $10^{z_{0}}1$ are subwords of $\mathbf{w}$ and thus, by (2.1), the inequality (4.2) holds for $n=0$, which is a contradiction. It follows that $z_{0}>z_{n}+1$ and consequently $z_{0}\geq z_{n}+3$ for all $n\geq 1$. Then $0110^{z_{n}}1$ and $10^{z_{n}+3}$ are subwords of $\mathbf{w}$. Thus, by (2.1), we have
$r+r^{2}-r^{3}+r^{z_{n}+4}-r^{z_{n}+4}(1-r)(1-r^{2})=(1+r^{z_{n}+4})(B-A)\\ \geq t_{-r}(0110^{z_{n}}1)-t_{-r}(10^{z_{n}+3})=r+r^{2}-r^{3}+r^{z_{n}+4}.$
Contradiction.
This finishes the proof of the claim $\mathbf{w}\in\mathcal{P}$. It follows that there exists an infinite word $\mathbf{s}$ such that $\mathbf{w}=D\mathbf{s}$ or $\mathbf{w}=T(D\mathbf{s})$. Assume that $\mathbf{s}\notin\mathcal{S}$. Then, by Lemma 4.2, there exists a finite word $\mathbf{u}$ such that $01\mathbf{u}1$ and $10\mathbf{u}0$ are subwords of $\mathbf{s}$. It follows that $011(D\mathbf{u})1$ and $100(D\mathbf{u})0$ are subwords of $\mathbf{w}$. Let $k$ be the length of $\mathbf{u}$. Then, by (2.1), we have
$r+r^{2}-r^{3}+r^{2k+4}-r^{2k+4}(1-r)(1-r^{2})=(1+r^{2k+4})(B-A)\\ \geq t_{-r}(011(D\mathbf{u})1)-t_{-r}(100(D\mathbf{u})0)=r+r^{2}-r^{3}+r^{2k+4}.$
Contradiction. Therefore, $\mathbf{s}\in\mathcal{S}$ and consequently $\mathbf{w}\in\mathcal{D}$.
For the other statements of the theorem, we may again assume $g=0$. Let $\mathbf{w}\in\mathcal{D}$ and let $\theta$ be the slope of $\mathbf{w}$. Let $W=\{t_{-r}(T^{n}\mathbf{w})\mid n\geq 0\}$. By Lemmas 4.9 and 4.10, since $T^{n}\mathbf{w}\in\mathcal{D}$ for all $n\geq 0$, $W$ is contained in the interval $t_{-r}(100D\mathbf{c}_{\theta}),t_{-r}(011D\mathbf{c}_{\theta})]$ of length $r+r^{2}-r^{3}$. By [LZ, Theorem 1.2], $W$ is not contained in any shorter interval. By Lemma 4.10, $t_{-r}(100D\mathbf{c}_{\theta})\in W$ if and only if $T^{n}\mathbf{w}=100D\mathbf{c}_{\theta}$ for some $n\geq 0$, and $t_{-r}(011D\mathbf{c}_{\theta})\in W$ if and only if $T^{m}\mathbf{w}=011D\mathbf{c}_{\theta}$ for some $m\geq 0$. If both conditions hold, then $T^{m+3}\mathbf{w}=D\mathbf{c}_{\theta}=T^{n+3}\mathbf{w}$ for $m\neq n$, which contradicts the aperiodicity of $\mathbf{w}$. Thus $W$ is contained in $t_{-r}(100D\mathbf{c}_{\theta}),t_{-r}(011D\mathbf{c}_{\theta}))$ or $(t_{-r}(100D\mathbf{c}_{\theta}),t_{-r}(011D\mathbf{c}_{\theta})]$. Furthermore, $W$ is contained in $(t_{-r}(100D\mathbf{c}_{\theta}),t_{-r}(011D\mathbf{c}_{\theta}))$ if and only if $\mathbf{w}\notin\mathcal{C}$. ∎
To show the transcendence in Theorem 1.2, we use the theorem of Ferenczi and Mauduit [FM, Proposition 2] as follows.
Lemma 4.11.
Let $b\geq 2$ be an integer and let $\mathbf{w}\in\mathcal{D}$. Then $t_{-1/b}(\mathbf{w})$ is transcendental.
Proof.
We have $\mathbf{w}=a^{l}D\mathbf{s}$, where $a\in\{0,1\}$, $l\geq 0$, and $\mathbf{s}$ is Sturmian. Then $t_{-1/b}(\mathbf{w})=t_{-1/b}(a^{l})+(-b)^{-l}(-b+1)t_{1/b^{2}}(\mathbf{s})$. By [FM, Proposition 2], $t_{1/b^{2}}(\mathbf{s})$ is transcendental. Thus the same holds for $t_{-1/b}(\mathbf{w})$. ∎
Proof of Theorem 1.2.
We apply Theorem 1.4 to $r=1/b$ as follows. Assume that $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is contained in an interval $I\subseteq\mathbb{R}/\mathbb{Z}$ of length $b^{-1}+b^{-2}-b^{-3}<1$. Choose $\eta\in\mathbb{R}$ such that $\pi(\eta)\notin I$. Then $I$ lifts to a unique interval $\tilde{I}\subseteq(\eta,\eta+1)$. For $n\geq 0$, let $y_{n}=\{\xi(-b)^{n}-\eta\}+\eta\in(\eta,\eta+1)$ and $v_{n}=-by_{n}-y_{n+1}\in\mathbb{Z}\cap(-(b+1)(\eta+1),-(b+1)\eta)$. Since $\pi(y_{n})=\pi(\xi(-b)^{n})$, we have $y_{n}\in\tilde{I}$ for all $n\geq 0$. Let $\mathbf{v}=v_{0}v_{1}\dots$. Then $t_{-1/b}(T^{n}\mathbf{v})=\sum_{i=0}^{\infty}v_{n+i}(-b)^{-(i+1)}=y_{n}$. Since $y_{0}$ is irrational, $\mathbf{v}$ is aperiodic. Then, by Theorem 1.4, there exist $g\in\mathbb{Z}$ and $\mathbf{w}=w_{1}w_{2}\dotso\in\mathcal{D}$ such that $v_{i}=-g+w_{i}$ for all $i\geq 1$. Thus $\pi(\xi)=\pi(y_{0})=\pi(g/(b+1)+t_{-1/b}(\mathbf{w}))$. By Lemma 4.11, it follows that $\xi$ is transcendental. The other assertions of Theorem 1.2 follow easily from the corresponding assertions of Theorem 1.4 and the formula $\pi(t_{-1/b}(T^{n}\mathbf{v}))=\pi((-b)^{n}t_{-1/b}(\mathbf{v}))$. ∎
Proof of Corollary 1.3.
While this can be deduced from Theorem 1.2, we deduce it more directly from Theorem 1.4, in parallel with the proof of Corollary 3.2. Take $0<r<1/2$. If there exists $\mathbf{u}$ such that $100\mathbf{u}\preceq_{\mathrm{alt}}T^{n}\mathbf{w}\preceq_{\mathrm{alt}}011\mathbf{u}$, then, by Lemma 4.7, $(t_{-r}(T^{n}\mathbf{w}))_{n\geq 0}$ is contained in the interval $t_{-r}(100\mathbf{u}),t_{-r}(011\mathbf{u})]$ of length $r+r^{2}-r^{3}$, which implies $\mathbf{w}\in\mathcal{D}$ by Theorem 1.2. Conversely, assume that $\mathbf{w}\in\mathcal{D}$. Let $\theta$ be the slope of $\mathbf{w}$. We have already proven $100D\mathbf{c}_{\theta}\preceq_{\mathrm{alt}}T^{n}\mathbf{w}\preceq_{\mathrm{alt}}011D\mathbf{c}_{\theta}$ in Lemma 4.9. If there exists an infinite word $\mathbf{v}$ over $\{0,1\}$ such that $100D\mathbf{c}_{\theta}\prec_{\mathrm{lex}}\mathbf{v}\preceq_{\mathrm{lex}}T^{n}\mathbf{w}$ for all $n\geq 0$, then, by Lemma 4.7, $(t_{-r}(T^{n}\mathbf{w}))_{n\geq 0}$ is contained in the interval $t_{-r}(\mathbf{v}),t_{-r}(011D\mathbf{c}_{\theta})]$ of length $r+r^{2}-r^{3}$, which contradicts [LZ, Theorem 1.2]. Thus $100D\mathbf{c}_{\theta}=\inf\nolimits^{\mathrm{alt}}_{n\geq 0}T^{n}\mathbf{w}$ and similarly $011D\mathbf{c}_{\theta}=\sup\nolimits^{\mathrm{alt}}_{n\geq 0}T^{n}\mathbf{w}$. ∎
Remark 4.12.
The interval with endpoints (1.3) in Theorem 1.2 is a neighborhood of $\pi([(g-1)/(b+1)-(b^{-2}-b^{-3}),g/(b+1)+(b^{-2}-b^{-3})])$, where $g\in\mathbb{Z}$. In particular, the union of all the open intervals of $\mathbb{R}/\mathbb{Z}$ of length $b^{-1}+b^{-2}-b^{-3}$ containing $(\pi(\xi(-b)^{n}))_{n\geq 0}$ for some irrational real number $\xi$ is the entire $\mathbb{R}/\mathbb{Z}$. This stands in contrast to the case of positive common ratio ([BD, Theorem 2.1] or Corollary 3.3), where the closure of each interval of length $1/b$ containing $(\pi(\xi b^{n}))_{n\geq 0}$ for some irrational real number $\xi$ is contained in $\pi((g/(b-1),(g+1)/(b-1)))$ for some integer $g$.
Dubickas ([Dubickas, Theorem 3], [Dubickas-neg, Theorem 2]) showed that for any $\epsilon>0$ and any irrational real number $\xi$, $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is not contained in any of the following four intervals:
$\pi([-A+\epsilon,A-\epsilon]),\quad\pi([\epsilon,B-\epsilon]),\quad\pi([1/2-A^{\prime}+\epsilon,1/2+A^{\prime}-\epsilon]),\quad\pi([1-B+\epsilon,1-\epsilon]),$
where $B=1-\prod_{k=1}^{\infty}(1-b^{-(2^{k}+(-1)^{k-1})/3})$, $A=(1-(1-b^{-1})P)/2$, $P=\prod_{k=0}^{\infty}(1-b^{-2^{k}})$, $A^{\prime}=A$ if $b$ is odd and $A^{\prime}=(1-P)/2$ if $b$ is even. Theorem 1.2 implies that the interval with endpoints (1.3) is not contained in any of the four intervals.